The primary coil of a transformer contains 24 turns of wire. The secondary coil of the transformer contains 1200 turns of wire. 


The primary coil is connected to an alternating potential difference of 6.0 V. Calculate the potential difference across the secondary coil. 


1
Expert's answer
2022-03-27T10:55:54-0400

"k=\frac{U_2}{U_1}=\frac{N_2}{N_1},"

"U_2=U_1\frac{N_2}{N_1}=300~V."